### v1

@@ Abstract
In this paper, we study second-order algorithms for solving nonconvex-strongly concave minimax problems, which have attracted much attention in recent years in many fields, especially in machine learning. We propose a gradient norm regularized trust region (GRTR) algorithm to solve nonconvex-strongly concave minimax problems, where the objective function of the trust region subproblem in each iteration uses a regularized version of the Hessian matrix, and the regularization coefficient and the radius of the ball constraint are proportional to the square root of the gradient norm. The iteration complexity of the proposed GRTR algorithm to obtain an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point is proved to be upper bounded by $\tilde{\mathcal{O}}(\rho^{0.5}\kappa^{1.5}\epsilon^{-3/2})$, where $\rho$ and $\kappa$ are the Lipschitz constant of the Jacobian matrix and the condition number of the objective function respectively, which matches the best known iteration complexity of second-order methods for solving nonconvex-strongly concave minimax problems. We further propose a Levenberg-Marquardt algorithm with a gradient norm regularization coefficient and use the negative curvature direction to correct the iteration direction (LMNegCur), which does not need to solve the trust region subproblem at each iteration. We also prove that the LMNegCur algorithm achieves an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point within $\tilde{\mathcal{O}}(\rho^{0.5}\kappa^{1.5}\epsilon^{-3/2})$ number of iterations. Numerical results show the efficiency of both proposed algorithms.
@@
Trust region, Levenberg-Marquardt algorithm, nonconvex-strongly concave minimax problems, second-order algorithms
@@ 1 Introduction
Three main types of optimization algorithms have been developed to solve the minimax problem (1.1), including zero-order, first-order, and second-order optimization algorithms, which exploit the function value, gradient, and Hessian information of the objective function, respectively. Compared with zero-order and first-order optimization algorithms, second-order optimization algorithms have received more attention due to their faster convergence speed. In addition, second-order optimization algorithms can handle saddle points and local minima more effectively, thereby increasing the probability of finding the global optimal solution. In this paper, we focus on the second-order optimization algorithms for solving (1.1).
@@
However, there are relatively few studies on second-order algorithms for solving nonconvex minimax optimization problems (1.1). Existing second-order algorithms can be divided into two categories, i.e., cubic regularization type algorithms Luo2022FindingSS ; Chen2023ACR and trust region type algorithms Yao2024TwoTR . For the cubic regularization type algorithms, Luo et al. Luo2022FindingSS proposed a Minimax Cubic Newton (MCN) algorithm, which can find an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point with an iteration complexity of $\mathcal{O}(\epsilon^{-3/2})$. MCN performs accelerated gradient ascent method to update $y$, which is then used to estimate the gradient and Hessian of $P(x)=\max_{y\in\mathbb{R}^{m}}f(x,y)$ involved in the cubic regularization update for $x$. Chen et al. Chen2023ACR proposed a Cubic-LocalMinimax algorithm with the same iteration complexity as Luo2022FindingSS . For the trust region type algorithms, Yao and Xu Yao2024TwoTR first proposed a minimax trust region (MINIMAX-TR) algorithm, which can be viewed as an inexact trust region method with fixed trust region radius of $\mathcal{O}(\sqrt{\epsilon})$. Nevertheless, this strategy is relatively conservative, leading to poor numerical performance. To overcome this shortcoming, they further proposed a minimax trust region algorithm with contractions and expansions (MINIMAX-TRACE) algorithm. Both algorithms have the same iteration complexity result as that in Luo2022FindingSS ; Chen2023ACR . However, the MINIMAX-TRACE algorithm needs function information of $P(x)$ to calculate the ratio of actual-to-predicted reduction and additional assumption that the gradient of $P(x)$ is upper bounded by a constant.
@@ 1.1 Contributions
In this paper, we focus on developing a gradient norm regularized second-order algorithm to solve the nonconvex-strongly concave minimax optimization problems. We summarize our contributions as follows.
@@
We first propose a gradient norm regularized trust region (GRTR) algorithm to solve the nonconvex-strongly concave minimax optimization problems, where the objective function of the trust region subproblem in each iteration adopts a regularized version of the Hessian matrix, and the regularization coefficient and the radius of the ball constraint are proportional to the root mean square of the gradient norm. It can be shown that the iteration complexity of the GRTR algorithm to obtain an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point is bounded by $\tilde{\mathcal{O}}(\rho^{0.5}\kappa^{1.5}\epsilon^{-3/2})$, where $\tilde{\mathcal{O}}$ hides only absolute constants and logarithmic factors. Ignoring the logarithmic factor, this iteration complexity is consistent with the state-of-art existing second-order algorithms Luo2022FindingSS ; Chen2023ACR ; Yao2024TwoTR . In addition, compared with the MINIMAX-TR algorithm Yao2024TwoTR , our method does not need to always fix the trust region radius, but the trust region radius varies according to the gradient norm, which leads to better numerical performance. Compared with the MINIMAX-TRACE algorithm Yao2024TwoTR , we do not require additional function information and do not need to assume that the gradient has a uniform upper bound.
@@
We further propose a Levenberg-Marquardt algorithm with a gradient norm regularization coefficient and use the negative curvature direction to correct the iteration direction (LMNegCur), which does not need to solve the trust region subproblem at each iteration. The LMNegCur algorithm can also achieve an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point within $\tilde{\mathcal{O}}(\rho^{0.5}\kappa^{1.5}\epsilon^{-3/2})$ number of iterations.
@@ 1.2 Related Works
There are also some existing research results on first-order algorithms for solving nonconvex-strongly concave minimax problems Bot ; Huang ; Jin ; Lin2019 ; Lin2020 ; Lu ; Rafique . To obtain an $\epsilon$-first-order stationary point of $P(x)$, or the stationary points of $f(x,y)$, these algorithms can achieve an iteration complexity of $\tilde{\mathcal{O}}\left(\kappa_{y}^{2}\epsilon^{-2}\right)$, where $\kappa_{y}$ is the condition number of $f(x,\cdot)$. In addition, Zhang et al. Zhang2021 can improve the iteration complexity to $\mathcal{O}\left(\sqrt{\kappa_{y}}\epsilon^{-2}\right)$ by introducing a general acceleration framework.
@@
There are also some research results on the zeroth-order method for solving the black-box minimax problems, where we can only access function values. For the nonconvex-strongly concave setting, various algorithms have been proposed, eg., the ZO-Min-Max algorithm Liu , the zeroth-order gradient descent multi-step ascent algorithm Wang , and the accelerated zeroth-order momentum descent ascent algorithm Huang . Among these algorithms, Huang achieves the best known iteration complexity of $\tilde{\mathcal{O}}(\kappa_{y}^{4.5}\epsilon^{-3})$, where $\kappa_{y}$ is the condition number for $f(x,\cdot)$. For the nonconvex-concave setting, Xu et al. xu21zeroth proposed a zero-order alternating stochastic gradient projection (ZO-AGP) algorithm with aniteration complexity of $\mathcal{O}(\epsilon^{-4})$. For the nonconvex-nonconcave setting, Xu et al. xu22zeroth proposed a zeroth-order alternating gradient descent ascent (ZO-AGDA) algorithm and a zeroth-order variance reduced alternating gradient descent ascent (ZO-VRAGDA) algorithm under the deterministic and the stochastic setting, respectivelly. The iteration complexity of the ZO-AGDA algorithm and the ZO-VRAGDA algorithm have been proved to be $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-3})$, respectively.
@@
2 A gradient norm regularized trust region algorithm
In this section, we propose a gradient norm regularized trust region (GRTR) algorithm for solving (1.1). The proposed algorithm is inspired by the universal trust-region (UTR) method Jiang2023AUT , which is a trust region variant for solving $\min_{x\in\mathbb{R}^{n}}P(x)$. In particular, at each iteration, the UTR method solves the following gradient-regularization trust-region subproblem:
@@ 2 A gradient norm regularized trust region algorithm
where $\sigma_{t}$ and $r_{t}$ are iteration-dependent parameters, the objective function of the trust region subproblem uses a regularized version of the Hessian matrix, and the regularization coefficient and the radius of the ball constraint are proportional to the root mean square of the gradient norm. The proposed GRTR algorithm can be regarded as an inexact generalization version of UTR for solving (1.1) with $P(x):=\max_{y\in\mathbb{R}^{m}}f(x,y)$, which consists of the following two important algorithmic components at each iteration:
  • Gradient ascent update: for a given $x_{t}$, run $N_{t}$ iterations of gradient ascent step to obtain an approximated maximizer
@@
  • Trust region update: for given $(x_{t},y_{t})$, compute a direction $s_{t}$ such that it is a solution of the following trust-region subproblem:
@@
Algorithm 1 A gradient norm regularized trust region (GRTR) algorithm
  Step 1: Input $x_{0},y_{0},\eta_{y},\sigma,r,\epsilon,N_{t}$; Set $t=1$, $x_{1}=x_{0}$.
@@
   (a): Set $k=0$, $y_{k}^{t}=y_{t-1}$.
   (b): Update $y_{k}^{t}$:
$y_{k+1}^{t}=y_{k}^{t}+\eta_{y}\nabla_{y}f(x_{t},y_{k}^{t}).$ (2.2)
@@
  Step 4: Update $s_{t}$ by solving the following trust region subproblem:
@@
Note that in the trust region subproblem (2.3), the trust region radius varies with the gradient norm. When $\|g_{t}\|\leqslant\epsilon$, we fix the trust region radius to $\mathcal{O}(\sqrt{\epsilon})$. If we set $\sigma=0$ and fix the trust region radius to $\mathcal{O}(\sqrt{\epsilon})$, then the GRTR algorithm is equivalent to the MINIMAX-TR algorithm Yao2024TwoTR . Compared with the MINIMAX-TRACE algorithm Yao2024TwoTR , the GRTR algorithm does not need to calculate the function value of $P(x)$. In the next section, we will prove that the iteration complexity of the GRTR algorithm to solve (1.1) for an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point of $P(x)$ is Note that in the trust region subproblem (2.3), the trust region radius varies with the gradient norm. When $\|g_{t}\|\leqslant\epsilon$, we fix the trust region radius to $\mathcal{O}(\sqrt{\epsilon})$. If we set $\sigma=0$ and fix the trust region radius to $\mathcal{O}(\sqrt{\epsilon})$, then the GRTR algorithm is equivalent to the MINIMAX-TR algorithm Yao2024TwoTR . Compared with the MINIMAX-TRACE algorithm Yao2024TwoTR , the GRTR algorithm does not need to calculate the function value of $P(x)$. In the next section, we will prove the iteration complexity of the GRTR algorithm to solve (1.1) for an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point of $P(x)$ is $\tilde{\mathcal{O}}(\rho^{0.5}\kappa^{1.5}\epsilon^{-1.5})$.

### v2

@@ Abstract
In this paper, we study second-order algorithms for solving nonconvex-strongly concave minimax problems, which have attracted much attention in recent years in many fields, especially in machine learning. We propose a gradient norm regularized trust-region (GRTR) algorithm to solve nonconvex-strongly concave minimax problems, where the objective function of the trust-region subproblem in each iteration uses a regularized version of the Hessian matrix, and the regularization coefficient and the radius of the ball constraint are proportional to the square root of the gradient norm. The iteration complexity of the proposed GRTR algorithm to obtain an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point is proved to be upper bounded by $\tilde{\mathcal{O}}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5})$, where $\mu$ is the strong concave coefficient, $\ell$ and $\rho$ are the Lipschitz constant of the gradient and Jacobian matrix respectively, which matches the best known iteration complexity of second-order methods for solving nonconvex-strongly concave minimax problems. We further propose a Levenberg-Marquardt algorithm with a gradient norm regularization coefficient and use the negative curvature direction to correct the iteration direction (LMNegCur), which does not need to solve the trust-region subproblem at each iteration. We also prove that the LMNegCur algorithm achieves an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point within $\tilde{\mathcal{O}}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5})$ number of iterations. The inexact variants of both algorithms can still obtain $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary points with high probability, but only require $\tilde{\mathcal{O}}(\ell^{2.25}\rho^{0.25}\mu^{-1.75}\epsilon^{-1.75})$ Hessian-vector products and $\tilde{\mathcal{O}}(\ell^{2}\rho^{0.5}\mu^{-2}\epsilon^{-1.5})$ gradient ascent steps. Numerical results show the efficiency of both proposed algorithms.
@@
trust-region, Levenberg-Marquardt algorithm, nonconvex-strongly concave minimax problems, second-order algorithms
@@ 1 Introduction
Three main types of optimization algorithms have been developed to solve the minimax problem (1.1), including zeroth-order, first-order, and second-order optimization algorithms, which exploit the function value, gradient, and Hessian information of the objective function, respectively. Compared with zeroth-order and first-order optimization algorithms, second-order optimization algorithms have received more attention due to their faster convergence speed. In addition, second-order optimization algorithms can handle saddle points and local minima more effectively, thereby increasing the probability of finding the global optimal solution. In this paper, we focus on the second-order optimization algorithms for solving (1.1).
@@
However, there are relatively few studies on second-order algorithms for solving nonconvex minimax optimization problems (1.1). Existing second-order algorithms can be divided into two categories, i.e., cubic regularization type algorithms Luo2022FindingSS ; Chen2023ACR and trust-region type algorithms Yao2024TwoTR . For the cubic regularization type algorithms, Luo et al. Luo2022FindingSS proposed a Minimax Cubic Newton (MCN) algorithm, which can find an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point with an iteration complexity of $\mathcal{O}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5})$. MCN performs accelerated gradient ascent method to update $y$, which is then used to estimate the gradient and Hessian of $P(x)=\max_{y\in\mathbb{R}^{m}}f(x,y)$ involved in the cubic regularization update for $x$. The proposed inexact variant of the MCN algorithm (IMCN) solves the cubic subproblem via perturbative gradient descent and matrix Chebyshev expansion, which requires a total of $\tilde{\mathcal{O}}(\ell^{2.5}\mu^{-1.5}\epsilon^{-2})$ Hessian-vector product calls and $\tilde{\mathcal{O}}(\ell^{2}\rho^{0.5}\mu^{-2}\epsilon^{-1.5})$ gradient ascent step calls to find an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point with high probability. Chen et al. Chen2023ACR proposed an inexact Cubic-LocalMinimax (ICLM) algorithm with the same complexity bound as Luo2022FindingSS . In addition to the cubic regularization type algorithm, the trust region type method is also a very representative second-order method, due to its robustness and excellent numerical performance in dealing with nonconvex problems. For the trust-region type algorithms, Yao and Xu Yao2024TwoTR first proposed a minimax trust-region (MINIMAX-TR) algorithm, which can be viewed as an inexact trust-region method with fixed trust-region radius of $\mathcal{O}(\sqrt{\epsilon})$, and has the same iteration complexity results as Luo2022FindingSS ; Chen2023ACR . Nevertheless, this strategy is relatively conservative, leading to poor numerical performance. To overcome this shortcoming, they further proposed a minimax trust-region algorithm with contractions and expansions (MINIMAX-TRACE) algorithm. However, the iteration complexity of the MINIMAX-TRACE algorithm is slightly worse, $\mathcal{O}(\ell^{4.5}\rho^{1.5}\mu^{-4.5}\epsilon^{-1.5})$, and needs function information of $P(x)$ to calculate the ratio of actual-to-predicted reduction and additional assumption that the gradient of $P(x)$ is upper bounded by a constant. There is also no analysis of the complexity of the Hessian-vector product that allows for inexact trust region subproblem solutions.
@@ 1.1 Contributions
In this paper, we focus on developing a more efficient and robust second-order algorithm to solve the nonconvex-strongly concave minimax optimization problems. We summarize our contributions as follows.
@@
We first propose a gradient norm regularized trust-region (GRTR) algorithm to solve the nonconvex-strongly concave minimax optimization problems, where the objective function of the trust-region subproblem in each iteration adopts a regularized version of the Hessian matrix, and the regularization coefficient and the radius of the ball constraint are proportional to the root mean square of the gradient norm. It can be shown that the iteration complexity of the GRTR algorithm to obtain an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point is bounded by $\tilde{\mathcal{O}}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5})$, where $\tilde{\mathcal{O}}$ hides only absolute constants and logarithmic factors. Ignoring the logarithmic factor, this iteration complexity is consistent with the state-of-art existing second-order algorithms Luo2022FindingSS ; Chen2023ACR ; Yao2024TwoTR . In addition, compared with the MINIMAX-TR algorithm Yao2024TwoTR , our method does not need to always fix the trust-region radius, but the trust-region radius varies according to the gradient norm, which leads to better numerical performance. Compared with the MINIMAX-TRACE algorithm Yao2024TwoTR , we do not require additional function information and do not need to assume that the gradient has a uniform upper bound.
To solve large-scale problems, we propose an inexact GRTR algorithm variant (IGRTR) that uses either truncated conjugate gradient or negative curvature to solve the trust region subproblem inexactly. The IGRTR algorithm requires a total of $\tilde{\mathcal{O}}(\ell^{2.25}\rho^{0.25}\mu^{-1.75}\epsilon^{-1.75})$ Hessian-vector products and $\tilde{\mathcal{O}}(\ell^{2}\rho^{0.5}\mu^{-2}\epsilon^{-1.5})$ gradient ascent steps to find $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary points with high probability. Compared with the IMCN algorithm Luo2022FindingSS and the ICLM algorithm Chen2023ACR , the proposed IGRTR algorithm has better complexity bounds in terms of Hessian-vector product calls, which is the main computational effort when solving large-scale problems.
@@
We further propose a Levenberg-Marquardt algorithm with a gradient norm regularization coefficient and use the negative curvature direction to correct the iteration direction (LMNegCur), which does not need to solve the trust-region subproblem at each iteration. The LMNegCur algorithm can also achieve an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point within $\tilde{\mathcal{O}}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5})$ number of iterations.
Table 1: Comparison of second-order algorithms for solving nonconvex-strongly concave minimax problems to $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary points.
Algorithm Type ##\## of outer iterations ##\## of Hessian-vector products
IMCN Luo2022FindingSS CR1 $\mathcal{O}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5})$ $\tilde{\mathcal{O}}(\ell^{2.5}\mu^{-1.5}\epsilon^{-2})$
ICLM Chen2023ACR CR $\mathcal{O}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5})$ $\tilde{\mathcal{O}}(\ell^{2.5}\mu^{-1.5}\epsilon^{-2})$
MINIMAX-TRYao2024TwoTR TR2 $\mathcal{O}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5})$ 4
MINIMAX-TRACEYao2024TwoTR TR $\mathcal{O}(\ell^{4.5}\rho^{1.5}\mu^{-4.5}\epsilon^{-1.5})$
IGRTR TR $\tilde{\mathcal{O}}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5})$ $\tilde{\mathcal{O}}(\ell^{2.25}\rho^{0.25}\mu^{-1.75}\epsilon^{-1.75})$
ILMNegCur LM3 $\tilde{\mathcal{O}}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5})$ $\tilde{\mathcal{O}}(\ell^{2.25}\rho^{0.25}\mu^{-1.75}\epsilon^{-1.75})$
  • 1
    “CR” denotes cubic regularization type algorithms.
  • 2
    “TR” denotes trust-region type algorithms.
  • 3
    “LM” denotes Levenberg-Marquardt type algorithms.
  • 4
    “–” indicates that the Hessian-vector product complexity of the proposed algorithm is not analyzed when the subproblems are not solved exactly.
@@ 1.2 Related Works
There are also some existing research results on first-order algorithms for solving nonconvex-strongly concave minimax problems Bot ; Jin ; Lin2019 ; Lin2020 ; Lu ; Rafique . To obtain an $\epsilon$-first-order stationary point of $P(x)$, or the stationary points of $f(x,y)$, these algorithms can achieve an iteration complexity of $\tilde{\mathcal{O}}(\kappa_{y}^{2}\epsilon^{-2})$, where $\kappa_{y}$ is the condition number of $f(x,\cdot)$. In addition, Zhang et al. Zhang2021 can improve the iteration complexity to $\tilde{\mathcal{O}}(\ell^{1.5}\mu^{-0.5}\epsilon^{-2})$ by introducing a general acceleration framework.
@@
There are also some research results on the zeroth-order method for solving the black-box minimax problems, where we can only access function values. For the nonconvex-strongly concave setting, various algorithms have been proposed, eg., the ZO-Min-Max algorithm Liu , the zeroth-order (stochastic) gradient descent ascent (ZO-GDA) algorithm Wang , the zeroth-order (stochastic) gradient descent multi-step ascent (ZO-GDMSA) algorithm Wang , and the zeroth-order variance reduced gradient descent ascent (ZO-VRGDA) algorithm Xu20 . Among these algorithms, the ZO-GDMSA algorithm achieves a complexity bound of $\tilde{\mathcal{O}}(\kappa_{y}^{2}\epsilon^{-2})$ in the deterministic setting, while the ZO-VRGDA algorithm achieves a complexity bound of $\mathcal{O}(\kappa_{y}^{3}\epsilon^{-3})$ in the stochastic setting, where $\kappa_{y}$ is the condition number for $f(x,\cdot)$. For the nonconvex-concave setting, Xu et al. xu21zeroth proposed a zeroth-order alternating stochastic gradient projection (ZO-AGP) algorithm with aniteration complexity of $\mathcal{O}(\epsilon^{-4})$. For the nonconvex-nonconcave setting, Xu et al. xu22zeroth proposed a zeroth-order alternating gradient descent ascent (ZO-AGDA) algorithm and a zeroth-order variance reduced alternating gradient descent ascent (ZO-VRAGDA) algorithm under the deterministic and the stochastic setting, respectivelly. The iteration complexity of the ZO-AGDA algorithm and the ZO-VRAGDA algorithm have been proved to be $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-3})$, respectively.
@@
2 A gradient norm regularized trust-region algorithm
In this section, we propose a gradient norm regularized trust-region (GRTR) algorithm for solving (1.1). The proposed algorithm is inspired by the universal trust-region (UTR) method Jiang2023AUT , which is a trust-region variant for solving $\min_{x\in\mathbb{R}^{n}}P(x)$. In particular, at each iteration, the UTR method solves the following gradient-regularization trust-region subproblem:
@@ 2 A gradient norm regularized trust-region algorithm
where $\sigma_{t}$ and $r_{t}$ are iteration-dependent parameters, the objective function of the trust-region subproblem uses a regularized version of the Hessian matrix, and the regularization coefficient and the radius of the ball constraint are proportional to the root mean square of the gradient norm. The proposed GRTR algorithm can be regarded as an inexact generalization version of UTR for solving (1.1) with $P(x):=\max_{y\in\mathbb{R}^{m}}f(x,y)$, which consists of the following two important algorithmic components at each iteration:
  • Gradient ascent update: for a given $x_{t}$, run $N_{t}$ iterations of Nesterov’s accelerated Nesterov2018Lectures gradient ascent step to obtain an approximated maximizer
@@
  • Trust-region update: for given $(x_{t},y_{t})$, compute a direction $s_{t}$ such that it is a solution of the following trust-region subproblem:
@@
Algorithm 1 A gradient norm regularized trust-region (GRTR) algorithm
  Step 1: Input $x_{0},y_{-1},\eta_{y},\theta,\sigma,r,N_{t},\epsilon$; Set $t=0$.
@@
   (a): Set $k=0$, $y_{k}^{t}=\tilde{y}_{k}^{t}=y_{t-1}$.
   (b): Update $y_{k}^{t}$ and $\tilde{y}_{k}^{t}$:
$y_{k+1}^{t}=\tilde{y}_{k}^{t}+\eta_{y}\nabla_{y}f(x_{t},\tilde{y}_{k}^{t}),~{}\tilde{y}_{k+1}^{t}=y_{k+1}^{t}+\theta(y_{k+1}^{t}-y_{k}^{t}).$ (2.2)
@@
  Step 4: Update $s_{t}$ by solving the following trust-region subproblem:
@@
Note that in the trust-region subproblem (2.3), the trust-region radius varies with the gradient norm. When $\|g_{t}\|\leqslant\epsilon$, we fix the trust-region radius to $\mathcal{O}(\sqrt{\epsilon})$. If we set $\sigma=0$ and fix the trust-region radius to $\mathcal{O}(\sqrt{\epsilon})$, then the GRTR algorithm is equivalent to the MINIMAX-TR algorithm Yao2024TwoTR . Compared with the MINIMAX-TRACE algorithm Yao2024TwoTR , the GRTR algorithm does not need to calculate the function value of $P(x)$. Note that in the trust-region subproblem (2.3), the trust-region radius varies with the gradient norm. When $\|g_{t}\|\leqslant\epsilon$, we fix the trust-region radius to $\mathcal{O}(\sqrt{\epsilon})$. If we set $\sigma=0$ and fix the trust-region radius to $\mathcal{O}(\sqrt{\epsilon})$, then the GRTR algorithm is equivalent to the MINIMAX-TR algorithm Yao2024TwoTR . Compared with the MINIMAX-TRACE algorithm Yao2024TwoTR , the GRTR algorithm does not need to calculate the function value of $P(x)$.
In the next section, we will prove that for solving (1.1) to obtain an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point of $P(x)$, the iteration complexity of the GRTR algorithm is $\tilde{\mathcal{O}}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5})$. Our complexity analysis extends the framework of monotonic function value reduction of $P(x)$ from Jiang2023AUT , categorizing iterations into distinct cases where each iteration either achieves sufficient descent in the objective value or induces linear contraction of the gradient norm. The key distinction lies in addressing the challenge of GRTR’s trust-region subproblem: the inherent inexactness from approximate gradient $g_{t}=\nabla_{x}f(x_{t},y_{t})$ and Hessian $H_{t}=H(x_{t},y_{t})$ introduces significant challenges absent in standard trust-region analyses. Our main contribution centers on rigorously quantifying these approximation errors through novel inner-loop convergence characterizations, coupled with systematic parameter selection strategies. This enables us to establish a noise-resistant iteration process with guaranteed monotonic descent while maintaining the overall iteration complexity - critical advancements that distinguish our analysis from conventional trust-region theory.